OpenStudy (anonymous):

If g(x) = x2 - 1, find g(3). A. 9 B. 5 C. 8 D. -1

OpenStudy (chriss):

C

hero (hero):

g(x) = x^2 - 1 g(3) = 3^2 - 1 g(3) = 9 - 1 g(3) = 8

OpenStudy (anonymous):

Chris is correct - but iomega, you should at try to post your guess -

hero (hero):

Also, try using the correct mathematical notation for exponent.

OpenStudy (anonymous):

it's exponent? i thought it is 2 times x

hero (hero):

Yes, it is exponent

OpenStudy (anonymous):

Yeah, sorry. I didn't notice. It is an exponent.